OpenStudy (anonymous):

FREE MEDAL WITH BEST ANSWER John and Tim are looking at the equation the square root of the quantity of 3 times x plus 4 equals square root of x . John says that the solution is extraneous. Tim says the solution is non-extraneous. Is John correct? Is Tim correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (anonymous):

@SBurchette please take a look?

OpenStudy (anonymous):

@TammisaurusRex take a look?

OpenStudy (sburchette):

So, we have that \[\sqrt{3x+4}=\sqrt{x}\] If we square both sides, we have \[3x+4=x\]Solving for x, we get x=-2. Substituting into the original equation, we get \[\sqrt{-2}=\sqrt{-2}\] Since the solution gives us a nonreal number, we conclude that the solution is extraneous.

OpenStudy (anonymous):

Thank you so much! I'm glad you explained. Now I'm awarding you a medal hehe. :)

OpenStudy (anonymous):

Thanks this helped alot